# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v3

@@ 5.7 Comparison, Sampling, and On-Chain Adjudication
Comparison and On-Chain Commitment. After all designated verifiers have posted their commitments (or the commitment deadline expires), the Scheduler distributes, off-chain, the inferencer-produced target hidden states $\{S_{i}^{(t)}\}_{t=1}^{T}$ for the relevant segment $i$ to each verifier assigned to that segment. Each verifier $v$ recomputes its local segment to obtain $\{\hat{S}_{i}^{(t)}\}_{t=1}^{T}$ and computes elementwise discrepancies $\Delta^{(t)}=\hat{S}_{i}^{(t)}-S_{i}^{(t)}$. Let $\phi(\cdot)$ denote a scalar aggregation function such as mean absolute error, and $\Psi(\cdot)$ a statistic (mean or clipped mean) summarizing deviations across tokens. The resulting score $T_{v}=\Psi(\{\phi(\Delta^{(t)})\}_{t=1}^{T})$ is compared against a tolerance $\tau$ calibrated for floating-point nondeterminism [rajbhandari2022deep]. The verifier’s binary verdict is
$b=\mathbb{I}[\,T_{v}\leq\tau\,]\in\{0,1\}.$ (26)
To prevent adaptive or copy-cat reporting, the verifier posts a hiding, binding commitment
$C=\textsf{Commit}(b,r),$ (27)
@@
where $r$ is fresh randomness; $C$ is posted during the commit phase, and $(b,r)$ is revealed later [bonneau2015sok, zhang2020delphi].
@@
Data Sampling. To ensure verifiers actually perform computation, the contract performs on-chain spot checks of sampled tensor elements [micali1999verifiable, chen2024zkml]. Once verdict commitments are posted, the Scheduler uses a verifiable random function (VRF) to generate unbiased sample indices:
$(r_{i},\pi_{i})$ $\leftarrow\textsf{VRF.Eval}(sk_{\text{sch}},\langle h,i,\{\gamma_{j}\}_{j=1}^{m}\rangle),$ (28)
$\mathcal{T}_{i}$ $\leftarrow\textsf{Sample}(\textsf{HKDF}(r_{i});n,q),$ (29)
@@
$\big\langle r_{i},\pi_{i},\mathcal{T}_{i},\mathbf{V}^{(T)}_{\mathcal{T}_{i}},\gamma^{(T)}_{i,\text{inf}},\texttt{Sig}^{(T)}_{i,\text{inf}},\{\texttt{MP}^{\text{inf}}_{j}\}_{j\in\mathcal{T}_{i}}\big\rangle,$ (30)
@@
    $\textsf{VRF.Verify}(vk_{\text{sch}},\langle h,i,\{\gamma_{j}\}\rangle,r_{i},\pi_{i})=1;$ (31)
  2. 2.
@@
    $\textsf{VerifySig}(\texttt{pk}_{i,\text{inf}},\gamma^{(T)}_{i,\text{inf}},\texttt{Sig}^{(T)}_{i,\text{inf}})=1;$ (32)
  3. 3.
@@
    $\forall j\!\in\!\mathcal{T}_{i},\;\textsf{MerkleVerify}(\gamma^{(T)}_{i,\text{inf}},\mathbf{V}^{(T)}_{j},\texttt{MP}^{\text{inf}}_{j})=1.$ (33)
These checks guarantee unbiased sampling, authentic roots, and valid inclusion proofs [merkle1989certified]. Any mismatch can be publicly demonstrated with the Scheduler’s relay signatures, triggering immediate slashing.
@@
(1) Commitment opening:
$\textsf{OpenCommit}(C;b,r)=1;$ (34)
@@
$\forall j\!\in\!\mathcal{T}_{i},\;\textsf{MerkleVerify}(\gamma^{(T)}_{i,\text{ver}},v_{j},\texttt{MP}^{\text{ver}}_{j})=1;$ (35)
(3) Value matching:
$\Delta_{j}=|v_{j}-\mathbf{V}^{(T)}_{j}|\leq\varepsilon,$ (36)
where $\varepsilon$ captures tolerated numeric noise [rajbhandari2022deep]. The agreement rate
$\hat{p}=\frac{1}{|\mathcal{T}_{i}|}\sum_{j\in\mathcal{T}_{i}}\mathbb{I}[\Delta_{j}\leq\varepsilon]$ (37)
must satisfy $\hat{p}\geq\theta$ for a positive verdict $b=1$.
Tail verifiers ($i=L$) perform an additional token-consistency check [brown2020language]:
$\textsf{TokenMatch}\big(\arg\max f_{\text{LM}}(S^{(T)}_{L+1}),y_{T}\big)=1,$ (38)
@@
Detection Probability and Parameterization. Let $\alpha$ denote the per-sample error rate when computation is incorrect. With $n$ sampled entries,
$P_{\text{pass}}\leq(1-\alpha)^{n},$ (39)
so to bound undetected cheating by $\delta$,
$n\geq\frac{1}{\alpha}\ln\frac{1}{\delta}.$ (40)
Calibration experiments provide empirical lower bounds on $\alpha$ and guide the choice of $\varepsilon$ and $\theta$, balancing tolerance and sensitivity [sun2024zkllm].
On-Chain Adjudication. After reveal, the contract aggregates revealed verdicts $\{b_{v}\}_{v\in\mathbf{V}}$ from $m$ verifiers. If a majority agree ($B\!\geq\!m/2$), the inference is accepted and honest verifiers share rewards. If most disagree but maintain internal consistency (pairwise hidden-state deviation $\leq\varepsilon$ over at least a $\tau$-fraction of pairs), the negative verdict stands and consistent verifiers are rewarded. Otherwise, the outcome is indeterminate; any verifier may later submit a succinct zero-knowledge proof demonstrating an error, triggering inferencer slashing and bounty payout [qu2025zkgpt, chen2024zkml].
@@
Dispute Resolution. The above mechanism guarantees correctness under a one-honest-verifier assumption but also supports escalation when committee capture is suspected. Because each verification task samples only a small committee ($m\ll N$), an adversary controlling a modest network fraction may occasionally dominate a sampled group. To mitigate this risk, if the first-round verdict is negative, the inferencer may request a reverification with an expanded committee of size $m^{\prime}>m$. If the second-round majority produces a positive outcome, the contract rescinds the inferencer’s penalty and slashes all first-round verifiers who voted False. Conversely, if the extended committee upholds the negative verdict, the inferencer bears the cost of the additional round.
This multi-stage design achieves adaptive robustness: typical cases finalize in one verification round with minimal on-chain load, while disputed cases escalate probabilistically until resolved with high confidence [bonneau2015sok, chen2024zkml].
6 Security Analysis
6.1 Threat Model and Assumptions
VeriLLM operates in a decentralized and adversarial environment where participants may deviate arbitrarily from the protocol to reduce computation cost or bias outcomes. We consider three potentially malicious roles—Inferencers, Verifiers, and a partially trusted or even Byzantine Scheduler. Smart contracts execute deterministically as deployed and are assumed correct by construction [bonneau2015sok]. The security goal is to guarantee inference correctness and accountability under these adversarial conditions.
@@ 6.1 Threat Model and Assumptions
  • EUF-CMA security of digital signatures. Node signatures are existentially unforgeable under chosen-message attack, preventing impersonation or undetected alteration of signed artifacts [micali1999verifiable].
  • Collision resistance of hash functions and binding of Merkle commitments. Once a node publishes a Merkle root $\gamma$, no two distinct tensors can produce the same root, ensuring immutability of hidden-state logs [merkle1989certified].
  • Unbiased verifiable randomness. VRF outputs used for role assignment and sampling are unpredictable prior to evaluation and publicly verifiable afterward, precluding adaptive bias [micali1999verifiable].
  • Soundness of zero-knowledge proofs. When zero-knowledge proofs attest signature validity or batched inclusion proofs, any false statement is rejected except with negligible probability [chen2024zkml, sun2024zkllm].
@@
  • One-honest-verifier assumption. At least one verifier per task is honest and online. This weaker-than-majority assumption suffices to guarantee safety via full re-execution or sampled audit detection [zhang2020delphi].
  • Eventual synchrony. The network eventually delivers all messages. Smart contracts enforce timeouts and a commit–then–reveal order to prevent adaptive inference of partial states.
  • Bounded numeric nondeterminism. Hardware and kernel-induced floating-point variance is modeled as small zero-mean noise within threshold $\varepsilon$; honest executions pass with overwhelming probability [rajbhandari2022deep].
These assumptions align with the threat model of verifiable ML systems such as zkML [chen2024zkml], zkLLM [sun2024zkllm], and decentralized inference frameworks [arun2025verde, mei2024fedmoe].
6.2 Security Objectives
VeriLLM provides cryptographically auditable correctness guarantees for large-scale LLM inference with minimal overhead. When all nodes behave honestly, both the off-chain and on-chain verification phases accept except with negligible probability due to benign numeric noise. If an Inferencer produces an incorrect result—by pruning layers, quantizing weights, truncating decoding, or recycling stale states—then one of three independent mechanisms detects it: (i) full hidden-state recomputation, (ii) commit–then–sample audit enforcing random openings, or (iii) a zero-knowledge proof of inconsistency from an honest verifier [qu2025zkgpt]. Thus, any materially incorrect inference is provably detectable on chain.
@@ 6.2 Security Objectives
Unpredictable sampling. Sampling indices for audits are derived from VRF outputs computed after all commitments are posted, eliminating adaptivity and ensuring independence between commitments and revealed indices [micali1999verifiable].
@@
6.3 Commit–Then–Sample: Binding and Unpredictability
The commit–then–sample protocol underpins VeriLLM’s verifiability. Each node first commits to its hidden-state tensor via Merkle root $r_{i}$; after all roots are posted, a public seed
$R=\mathsf{Hash}(r_{1},r_{2},\dots,r_{n})$ (41)
is computed to derive sampling positions. This sequencing guarantees both binding (immutability) and unpredictability (fair randomness).
Lemma 1 (Binding).
@@ 6.3 Commit–Then–Sample: Binding and Unpredictability
Together, these properties ensure that a node committing to falsified or incomplete states cannot anticipate future audits and will be detected with overwhelming probability. This commit–then–sample discipline yields deterministic accountability and makes deviation economically irrational in expectation [zhang2020delphi, chen2024zkml].
6.4 Analysis of Concrete Attacks

### v4

@@ 5.7. Comparison, Sampling, and On-Chain Adjudication
Comparison and On-Chain Commitment. After all designated verifiers have posted their commitments (or the commitment deadline expires), the Scheduler distributes, off-chain, the inferencer-produced target hidden states $\{S_{i}^{(t)}\}_{t=1}^{T}$ for the relevant segment $i$ to each verifier assigned to that segment. Each verifier $v$ recomputes its local segment to obtain $\{\hat{S}_{i}^{(t)}\}_{t=1}^{T}$ and computes elementwise discrepancies $\Delta^{(t)}=\hat{S}_{i}^{(t)}-S_{i}^{(t)}$. Let $\phi(\cdot)$ denote a scalar aggregation function such as mean absolute error, and $\Psi(\cdot)$ a statistic (mean or clipped mean) summarizing deviations across tokens. The resulting score $T_{v}=\Psi(\{\phi(\Delta^{(t)})\}_{t=1}^{T})$ is compared against a tolerance $\tau$ calibrated for floating-point nondeterminism (Rajbhandari et al., 2022). The verifier’s binary verdict is
(21) $b=\mathbb{I}[\,T_{v}\leq\tau\,]\in\{0,1\}.$
@@
To prevent adaptive or copy-cat reporting, the verifier posts a hiding, binding commitment $C=\textsf{Commit}(b,r)$, where $r$ is fresh randomness; $C$ is posted during the commit phase, and $(b,r)$ is revealed later (Bonneau et al., 2015; Zhang et al., 2020).
@@
Data Sampling. To ensure verifiers actually perform computation, the contract performs on-chain spot checks of sampled tensor elements (Micali et al., 1999; Chen et al., 2024). Once verdict commitments are posted, the Scheduler uses a verifiable random function (VRF) to generate unbiased sample indices:
(22) $(r_{i},\pi_{i})$ $\leftarrow\textsf{VRF.Eval}(sk_{\text{sch}},\langle h,i,\{\gamma_{j}\}_{j=1}^{m}\rangle),$
(23) $\mathcal{T}_{i}$ $\leftarrow\textsf{Sample}(\textsf{HKDF}(r_{i});n,q),$
@@
(24) $\big\langle r_{i},\pi_{i},\mathcal{T}_{i},\mathbf{V}^{(T)}_{\mathcal{T}_{i}},\gamma^{(T)}_{i,\text{inf}},\texttt{Sig}^{(T)}_{i,\text{inf}},\{\texttt{MP}^{\text{inf}}_{j}\}_{j\in\mathcal{T}_{i}}\big\rangle,$
@@
    (25) $\textsf{VRF.Verify}(vk_{\text{sch}},\langle h,i,\{\gamma_{j}\}\rangle,r_{i},\pi_{i})=1;$
  2. (2)
@@
    (26) $\textsf{VerifySig}(\texttt{pk}_{i,\text{inf}},\gamma^{(T)}_{i,\text{inf}},\texttt{Sig}^{(T)}_{i,\text{inf}})=1;$
  3. (3)
@@
    (27) $\forall j\!\in\!\mathcal{T}_{i},\;\textsf{MerkleVerify}(\gamma^{(T)}_{i,\text{inf}},\mathbf{V}^{(T)}_{j},\texttt{MP}^{\text{inf}}_{j})=1.$
These checks guarantee unbiased sampling, authentic roots, and valid inclusion proofs (Merkle, 1989). Any mismatch can be publicly demonstrated with the Scheduler’s relay signatures, triggering immediate slashing.
@@
(1) Commitment opening: $\textsf{OpenCommit}(C;b,r)=1$;
@@
(28) $\forall j\!\in\!\mathcal{T}_{i},\;\textsf{MerkleVerify}(\gamma^{(T)}_{i,\text{ver}},v_{j},\texttt{MP}^{\text{ver}}_{j})=1;$
(3) Value matching: $\Delta_{j}=|v_{j}-\mathbf{V}^{(T)}_{j}|\leq\varepsilon$, where $\varepsilon$ captures tolerated numeric noise (Rajbhandari et al., 2022). The agreement rate
(29) $\hat{p}=\frac{1}{|\mathcal{T}_{i}|}\sum_{j\in\mathcal{T}_{i}}\mathbb{I}[\Delta_{j}\leq\varepsilon]$
must satisfy $\hat{p}\geq\theta$ for a positive verdict $b=1$. Tail verifiers ($i=L$) perform an additional token-consistency check (Brown et al., 2020):
(30) $\textsf{TokenMatch}\big(\arg\max f_{\text{LM}}(S^{(T)}_{L+1}),y_{T}\big)=1,$
@@
Detection Probability and Parameterization. Let $\alpha$ denote the per-sample error rate when computation is incorrect. With $n$ sampled entries, $P_{\text{pass}}\leq(1-\alpha)^{n}$, so to bound undetected cheating by $\delta$, $n\geq\frac{1}{\alpha}\ln\frac{1}{\delta}$. Calibration experiments provide empirical lower bounds on $\alpha$ and guide the choice of $\varepsilon$ and $\theta$, balancing tolerance and sensitivity (Sun et al., 2024).
On-Chain Adjudication. After reveal, the contract aggregates revealed verdicts $\{b_{v}\}_{v\in\mathbf{V}}$ from $m$ verifiers. If a majority agree ($B\!\geq\!m/2$), the inference is accepted and honest verifiers share rewards. If most disagree but maintain internal consistency (pairwise hidden-state deviation $\leq\varepsilon$ over at least a $\tau$-fraction of pairs), the negative verdict stands and consistent verifiers are rewarded. Otherwise, the outcome is indeterminate; any verifier may later submit a succinct zero-knowledge proof demonstrating an error, triggering inferencer slashing and bounty payout (Qu et al., 2025b; Chen et al., 2024).
@@
Dispute Resolution. The above mechanism also supports escalation when committee capture is suspected. Because each verification task samples only a small committee ($m\ll N$), an adversary controlling a modest network fraction may occasionally dominate a sampled group. To mitigate this risk, if the first-round verdict is negative, the inferencer may request a reverification with an expanded committee of size $m^{\prime}>m$. If the second-round majority produces a positive outcome, the contract rescinds the inferencer’s penalty and slashes all first-round verifiers who voted False. Conversely, if the extended committee upholds the negative verdict, the inferencer bears the cost of the additional round.
This multi-stage design achieves adaptive robustness: typical cases finalize in one verification round with minimal on-chain load, while disputed cases escalate probabilistically until resolved with high confidence (Bonneau et al., 2015; Chen et al., 2024).
6. Security Analysis
6.1. Threat Model and Assumptions
VeriLLM operates in a decentralized and adversarial environment where participants may deviate arbitrarily from the protocol to reduce computation cost or bias outcomes. We consider three potentially malicious roles—Inferencers, Verifiers, and a partially trusted or even Byzantine Scheduler. Smart contracts execute deterministically as deployed and are assumed correct by construction (Bonneau et al., 2015). The security goal is to guarantee inference correctness and accountability under these adversarial conditions.
@@ 6.1. Threat Model and Assumptions
  • EUF-CMA security of digital signatures. Node signatures are existentially unforgeable under chosen-message attack, preventing impersonation or undetected alteration of signed artifacts (Micali et al., 1999).
  • Collision resistance of hash functions and binding of Merkle commitments. Once a node publishes a Merkle root $\gamma$, no two distinct tensors can produce the same root, ensuring immutability of hidden-state logs (Merkle, 1989).
  • Unbiased verifiable randomness. VRF outputs used for role assignment and sampling are unpredictable prior to evaluation and publicly verifiable afterward, precluding adaptive bias (Micali et al., 1999).
@@
  • Global honest majority. For each task, the pool of latent verifiers contains a majority of honest nodes during the verification process. This assumption is more robust than the traditional sample honest majority assumption, as even the committee contains a majority of dishonest nodes, incorrect execution can still be detected via the disputing and re-verification process.
  • Bounded numeric nondeterminism. Hardware and kernel-induced floating-point variance is modeled as small zero-mean noise within threshold $\varepsilon$; honest executions pass with overwhelming probability (Rajbhandari et al., 2022).
These assumptions align with the threat model of verifiable ML systems such as zkML (Chen et al., 2024), zkLLM (Sun et al., 2024), and decentralized inference frameworks (Arun et al., 2025; Mei et al., 2024).
6.2. Security Objectives
VeriLLM provides cryptographically auditable correctness guarantees for large-scale LLM inference with minimal overhead. When all nodes behave honestly, both the off-chain and on-chain verification phases accept except with negligible probability due to benign numeric noise. If an Inferencer produces an incorrect result—by pruning layers, quantizing weights, truncating decoding, or recycling stale states—then one of three independent mechanisms detects it: (i) full hidden-state recomputation, (ii) commit–then–sample audit enforcing random openings, or (iii) a zero-knowledge proof of inconsistency from an honest verifier (Qu et al., 2025b). Thus, any materially incorrect inference is provably detectable on chain.
@@ 6.2. Security Objectives
Unpredictable sampling. Sampling indices for audits are derived from VRF outputs computed after all commitments are posted, eliminating adaptivity and ensuring independence between commitments and revealed indices (Micali et al., 1999).
@@
6.3. Commit–Then–Sample: Binding and Unpredictability
The commit–then–sample protocol underpins VeriLLM’s verifiability. Each node first commits to its hidden-state tensor via Merkle root $r_{i}$; after all roots are posted, a public seed $R=\mathsf{Hash}(r_{1},r_{2},\dots,r_{n})$ is computed to derive sampling positions. This sequencing guarantees both binding (immutability) and unpredictability (fair randomness).
Lemma 6.1 (Binding).
@@ 6.3. Commit–Then–Sample: Binding and Unpredictability
Together, these properties ensure that a node committing to falsified or incomplete states cannot anticipate future audits and will be detected with overwhelming probability. This commit–then–sample discipline yields deterministic accountability and makes deviation economically irrational in expectation (Zhang et al., 2020; Chen et al., 2024).
6.4. Collusion, Censorship, and Dispute
Collusion among verifiers. A coalition may attempt to fabricate mutually consistent hidden states and corresponding Merkle roots such that on-chain openings match internally forged commitments. The commit–then–sample discipline prevents any post hoc adaptation: once a root is posted, the coalition must open VRF-chosen indices consistent with that fixed root. To ensure semantic correctness with respect to the prescribed model, VeriLLM introduces two safeguards: (i) token–state cross-checks at the tail segment, where verifiers decode from recomputed last-layer states and compare tokens to the published outputs; and (ii) a zero-knowledge escalation path, in which any single honest verifier can submit a succinct proof that a challenged prefix of the transcript violates the model’s forward mapping or decoding rule (Chen et al., 2024; Sun et al., 2024; Qu et al., 2025b). The contract verifies the proof on-chain, finalizing the inconsistency independently of the coalition’s reports.
Censorship resistance. All commitments are recorded on-chain, and verifier selection is derived from verifiable randomness. To suppress detection, an adversary must corrupt or block all verifiers selected for a given task. Under the one-honest-verifier assumption, there exists at least one verifier that remains honest and available, which suffices to enable offline recomputation, successful sampled audits, or the submission of a zero-knowledge proof of inconsistency. A complete denial-of-service attack is detectable via missed protocol deadlines, which trigger timeout-based remediation mechanisms.
7. Experiment and Evaluation
We perform comprehensive experiments on real-world prototype systems, and report the results in this section. The systematic implementation details are reported in Appendix D.
7.1. Main Results
We empirically measured the divergence of hidden states produced during inference across heterogeneous compute devices (tested on Qwen2.5-7B-Instruct (Team, 2024); additional models are forthcoming). Although floating-point non-determinism across GPU vendors introduces small numeric drift (Gupta et al., 2015; Wang et al., 2021), these deviations are limited in magnitude and statistically distinguishable. Consequently, the hidden-state comparison test can, with high probability, determine whether two traces originate from the same underlying model computation.
We further evaluated replacing full-precision models with quantized variants. Hidden states computed by quantized models exhibit substantially larger—and statistically significant—bias relative to the full-precision baseline, consistent with prior analyses of quantization-induced drift (Banner et al., 2018; Sun et al., 2024). Hence, simple statistical tests cleanly separate hidden states from normal inference versus those produced under precision degradation or adversarial manipulation.
Table 2. Comparison of Hidden-State Metrics for Qwen2.5-7B-Instruct Inference on M4 and RTX 5090.
# Exact Exp Match ($|\Delta|$) Exp Mismatch ($|\Delta|$) Mean $\epsilon$
>0.2>0.2 <0.2<0.2 >5>5 <5<5
64 4 13 48 2 1 0.009
256 23 46 201 7 2 0.006
1024 87 191 802 23 7 0.003
3584 1240 1 2343 0 1 0.002
Table 3. Comparison of Hidden-State Metrics for Qwen2.5-7B-Instruct-AWQ inference on RTX 5090 with verification on Mac M4 CPU using Qwen2.5-7B-Instruct.
# Exact Exp Match ($|\Delta|$) Exp Mismatch ($|\Delta|$) Mean $\epsilon$
>0.2>0.2 <0.2<0.2 >5>5 <5<5
64 0 39 13 7 5 -0.022
256 2 161 53 23 19 0.021
1024 7 653 201 99 71 -0.019
3584 29 2276 712 346 250 $0.014$
Bit-aware statistical comparison. Building on these results, we propose a feature-based comparison of hidden states. Given reference $S$ and candidate $\bar{S}$, each floating-point element is decomposed into sign, exponent, and mantissa (Wang et al., 2021). With tolerances $e_{w}>0$ (e.g., $e_{w}=10^{-2}$) and $e_{m}\!\in\!\mathbb{N}$ (e.g., 7 mantissa steps), we define:
(1) Exponent-first checks. $P_{e}$ — proportion of exponent mismatches; $P_{m}$ — fraction of large mantissa deviations among exponent mismatches; $P_{w}$ — fraction of small mantissa deviations among exponent matches. (2) Mean discrepancy. $e=\frac{1}{|\Delta|}\sum_{\delta\in\Delta}\delta$
Empirically calibrated acceptance thresholds are:
Off-chain: $P_{e}{<}0.05$, $P_{m}{>}0.75$, $P_{w}{>}0.80$, $e\in[-0.01,0.01]$
On-chain: $P_{e}{<}0.08$, $P_{m}{>}0.70$, $P_{w}{>}0.75$, $e\in[-0.02,0.02]$
These rules balance detection sensitivity with tolerance for benign noise, building on calibration methodologies from prior verifiable ML studies (Zhang et al., 2020; Chen et al., 2024). We validate these thresholds through comprehensive sensitivity analysis (Appendix F).
7.2. Performance and Overhead Breakdown
We evaluate VeriLLM across five dimensions: throughput and latency, on-chain cost, scalability, ablation, and failure injection (Bonneau et al., 2015). All tests use the 10-node setup described in Section VI.
7.2.1. Throughput and Latency
For a 7B-parameter model (512-token prompt, 128-token output), total latency per request is 1.37 s: inference 93.1 %, scheduler relay 3.4 %, and verification 3.5 %. Verification overhead averages 0.78 % of total inference time, consistent with lightweight cryptographic pipelines (Micali et al., 1999). Increasing verifier parallelism beyond three yields diminishing returns as GPU prefill already saturates compute throughput.
Table 4. Latency Breakdown for Qwen2.5-7B-Instruct (128-token output).
Component Mean (ms) Share (%) Cumulative (%)
Prefill + Decode (Inference) 1278 93.1 93.1
Scheduler Relay / Signatures 46 3.4 96.5
Verifier Prefill (Parallel) 33 2.4 98.9
On-chain Commit + Reveal 15 1.1 100.0
7.2.2. On-chain Cost Analysis
Each inference produces two transactions: (i) commit (108 k gas, 2.9 kB) and (ii) reveal (134 k gas, 3.2 kB). Merkle proofs require 8–10 hashes per sample, yielding $\approx$ 0.004 USD per verification at current Ethereum L2 prices (Chainlink Labs, 2023; Ethereum Foundation, 2024). VRF verification adds ¡ 4 % of total cost.
7.2.3. Scalability Experiments
Verification cost per token decreases inversely with verifier count $m$ until m>6m\!>\!6, where on-chain aggregation dominates. Prefill parallelization across $L$ segments achieves near-linear scaling $T_{\text{verify}}\!\approx\!T_{\text{prefill}}/L$, consistent with prior verifiable-inference frameworks (Arun et al., 2025).
7.2.4. Ablation Study
7.2.5. Failure Injection and Detection Latency
We emulate four attack classes—quantization, early termination, forged output, and lazy verification—via runtime perturbations (Zhang et al., 2020; Sun et al., 2024). Table 5 summarizes detection latency (from deviation to slash). All attacks are detected and finalized within 1–3 s, confirming that the commit–sample–reveal protocol yields rapid, tunable defense.
Table 5. Detection latency under injected attacks.
Attack Type Detection Probability Mean Detection Latency (s)
Quantization (W8A8) $99.9\%$ 1.7
Early Termination $100\%$ 1.2
Forged Output $99.8\%$ 2.4
Lazy Verification $99.9\%$ 2.8
8. Conclusion
This work presents VeriLLM, a publicly verifiable framework for decentralized large language model inference that achieves correctness and accountability with approximately 1% verification overhead. By combining Merkle-based commitments, VRF-driven sampling, and on-chain adjudication, VeriLLM allows any participant to audit inference without relying on a trusted majority or centralized coordinator. We formalize the protocol, analyze its security, and show that malicious deviations such as quantization, early termination, or forged outputs are detected with high probability, while cross-device floating-point noise remains bounded and statistically separable for lightweight hidden-state comparison. Experiments demonstrate efficient off-chain and on-chain checks that support scalable deployment across heterogeneous environments. By decoupling verification from replication and embedding transparency into inference traces, VeriLLM enables practical, trustworthy decentralized AI services; future directions include lightweight zero-knowledge proofs, support for multimodal models, and decentralized scheduling and incentive mechanisms.
Appendix A Analysis of Concrete Attacks